1) A cat of mass 1kg leaps to try to catch a bird. If the cat's jump was at an angle of 60 degrees with an initial velocity of 3 m/s, it's vertical displacement at the highest point of its trajectory is:

2) Suppose another cat of mass 2kg make the jump with the same initial velocity (3 m/s at an angle of 60 degrees with the horizontal). The total time this heavier cat is in the air is:

Please use a lot of details when explaining!
Answer #1


1) initial vertical velocity, u = 3 x Sin 60o = 2.6 m/s

at highest point vertical velocity, v = 0 m/s

acceleration = - 9.8 m/s2

using third equation

v2 - u2 = 2aS

S = v2 - u2 / 2a

substituting

S = 0 - 2.62 / 2 x -9.8 = 6.76 / 19.6 = 0.35 m

2) time of flight = 2 u Sin 60 / g = 2 x 2.6 /9.8 = 0.53 s